# Magnetic properties of monolayer, multilayer, and bulk CrTe2

### v2

@@ III Results
Figure 7: (Color online) Exchange interaction in bilayer CrTe2 at $\beta=18$ eV-1 (a) between the atoms of the same layer $J^{11}_{\mathbf{q}}$ and (b) between the atoms of different layers $J^{12}_{\mathbf{q}}$. Red solid line (squares) corresponds to DFT+DMFT approach in paramagnetic phase, blue dashed line (triangles) is the result of the DFT approach in FM phase. Full (open) symbols in (b) correspond to the real (imaginary) part.
We also used alternative method to obtain exchange interactions, based on comparison of energies of various spin configurations li_tunable_2021 , see Appendix B. The comparison of the obtained exchange interactions is presented in Table 1. All DFT approaches, except Ref. li_tunable_2021 yield the same signs of nearest- and next-nearest neighbor exchange interactions, with different magnitude. In particular, the nearest neighbour exchange interaction is negative (except Ref. li_tunable_2021 ) showing the tendency to antiferromagnetism. We find the sAFM-zz configuration in DFT approach 10 meV/f.u. lower than FM. At the same time, in DMFT approach we find weak ferromagnetic nearest neighbor exchange interaction, and the tendency to antiferromagnetism originates from the next to next nearest neighbors and further interactions.
To study effect of interlayer coupling we consider bilayer and trilayer CrTe2. The band structure of bilayer CrTe2 in paramagnetic phase is shown in Fig. 6. One can see that due to bilayer splitting the central sheet of the Fermi surface appears at the $\Gamma$ point. The electronic self-energies are shown in Fig. 2. One can see that the electronic damping for the bilayer system is even larger than for the single-layer one, and non-quasiparticle form of the self-energy is also more pronounced.
@@
We consider the interaction between the atoms of the same Cr plane (intralayer part), $J^{rr}_{\mathbf{q}}$ and the interplane interaction $J^{rr^{\prime}}_{\mathbf{q}}$ with $r\neq r^{\prime}$. From the momentum dependence of intralayer exchange interactions $J^{11}_{\mathbf{q}}=J^{22}_{\mathbf{q}}$ (see Fig. 7a) one can see that additional sheet of the Fermi surface strongly affects exchange interactions: the intralayer interaction becomes maximal near $\Gamma$ point of the Brillouin zone, while the height of the local maximum at the $K$ point decreases in comparison to the single-layer system. The obtained maximum of the intralayer exchange interaction can be contrasted to the results of DFT approach of the present study (shown by dashed line in Fig. 7a) and earlier DFT study zhu_insight_2023 , where the antiferromagnetic ground state was obtained. In case of bilayer system, both DFT and DFT+DMFT results for exchange interactions are weakly temperature dependent because of the absence of nesting and shift of the peak of the density of states from the Fermi level.
@@
Figure 10: (Color online) Real parts of the exchange interactions $J_{{ij},{\mathbf{q}}}$ in trilayer system at $\beta=18$ eV-1 in DFT+DMFT approach. The pairs of indexes $ij$ at the exchange interactions correspond to the pairs of layers (1-bottom, 2-middle, 3-top layer)
To demonstrate the possibility of intralayer ferromagnetic order in the bilayer CrTe2, we also show the temperature dependence of the inverse uniform intralayer susceptibility $(\chi^{11,22}_{{\mathbf{q}}=0})^{-1}$ in Fig. 8. One can see that in contrast to single-layer CrTe2 the inverse intralayer susceptibility vanishes at the DMFT Curie temperature $T_{C}^{\rm DMFT,{\text{2}{\text{-}}layer}}\simeq 400K$. Similarly to the single-layer system, this DMFT Curie temperature does not determine the true Curie temperature of the system, but rather the scale of the onset of strong ferromagnetic correlations. The temperature dependence of the local magnetic susceptibility (see the inset of Fig. 8) is similar to the single-layer case, and it is characterized by the magnetic moment $\mu^{2}_{\rm loc}=33.1\mu_{B}^{2}$, which is also close to that for the single-layer compound. Therefore the strong change of the non-local properties when passing from single- to bilayer CrTe2 almost does not affect the local properties.
@@
Let us consider what changes in DFT+DMFT results when adding one more layer in trilayer CrTe2. The Fermi surface possesses on additional sheet near the $\Gamma$ point (see Fig. 9). We find that the self-energies of trilayer system (not shown) are comparable to those of the bilayer one in Fig. 2. The maximum of the intralayer exchange interaction (see Fig. 10) shifts to the $\Gamma$ point (instead of being in the near vicinity of the $\Gamma$ point in bilayer system), which shows further stabilization of intralayer ferromagnetic order with increase of the number of layers. The exchange interaction between adjacent layers is negative (i.e. antiferromagnetic), similarly to the bilayer compound, with the maximum of the interlayer exchange interactions at the incommensurate positions in $\Gamma-M$ and $\Gamma-K$ directions.
@@
Figure 12: (Color online) Exchange interaction in bulk CrTe2
at $\beta=24$ eV-1. The notations are the same as in Fig. 5.
@@
To understand the effect of adding more layers in thick films, we consider bulk CrTe2. The obtained band structure and Fermi surface (Fig. 11) reproduce those of previous study freitas_ferromagnetism_2015 . The Fermi surface sheet closest to the $\Gamma$ point acquires the dispersion along $k_{z}$ axis in agreement with the results for the trilayer system. The exchange interactions at $\beta=24$ eV-1 (see Fig. 12) has local maximum at the $\Gamma$ point, but its value at the A point of the Brillouin zone is somewhat larger than at $\Gamma$, showing the tendency to alternating orientation between the adjacent planes, similarly to that obtained for 2- and 3-layer system. However, with reducing temperature (see Fig. 13), we observe a clear tendency of changing the global maximum of exchange interaction to $\Gamma$ point, which happens just at $T_{C}\simeq 250$ K, comparable to the experimental Curie temperature. This implies formation of low temperature FM phase (in agreement with the experimental data), which changes upon heating to the state with long- or short-range antiferromagnetic order. We expect the long range antiferromagnetic order above $T_{C}$ is destroyed by fluctuations. Therefore, we find an intriguing physical picture of competition of different types of magnetic correlations, which occurs just above the Curie temperature.
In view of the obtained results for the bilayer and trilayer systems, with decrease of the number of layers we expect the interlayer antiferromagnetic and incommensurate orders become progressively more favourable. Therefore, we expect the transition to interlayer incommensurate or antiferromagnetic states with decrease of the number of layers. In view of the delicate competition of various states, additional interactions, not included in the considered model (such as $d$-$p$ and $p$-$p$ interactions) may be important for obtaining critical number of layers for this transition, cf. Ref. Interlayer .
@@ IV Conclusions
In summary, we have investigated magnetic properties of monolayer, bilayer, trilayer, and bulk CrTe2. In monolayer system in both, DFT and DFT+DMFT approaches we find preference for the incommensurate magnetic order with the wave vector ${\mathbf{Q}}_{K}$, which corresponds to 120°  spin alignment. This order competes with AFM-zz magnetic order, earlier suggested for monolayer system.
In the bilayer and trilayer CrTe2 in DFT+DMFT approach we find, in contrast to the DFT results, the tendency to the intralayer ferromagnetic order, which appear due to electronic correlations. At the same time, similarly to previous DFT$+U$ studies we find the tendency to the interlayer AFM order. This tendency is preserved at not too low temperatures in bulk CrTe2, and changes to FM order at low temperatures.
Most important result of the present study is the possibility of stabilization of intralayer FM order in CrTe2 by correlations for number of layers bigger than one. Description of the transition from FM to AFM interlayer coupling with the decrease of the number of layers requires considering more sophisticated models, including non-local $d$-$p$ (as well as $p$-$p$) interaction. The FM order in single-layer CrTe2, observed experimentally zhang_room-temperature_2021 ; meng_anomalous_2021 ; liu_wafer-scale_2023 , requires further studies, and likely appears as an effect of substrate, cf. Ref. wang_strain-_2024 .
For future studies, account of full SU(2) Coulomb interaction, as well as the effect of the spin-orbit coupling would be desirable. Also, considering the effect of substrate, e.g., within the resently proposed approach of Ref. Substr is of certain interest.
@@ Acknowledgements
We are grateful I. A. Goremykin for the help with using Quantum Espresso and Wannier90 packages and discussions, and to A. K. Nukhov for discussions of the obtained results. The paramagnetic DFT+DMFT calculations were supported by the Russian Science Foundation (Project No. 24-12-00186).
Appendix A Single layer band structure in the ferromagnetic phase
In Fig. 14 we show band structure (cf. Refs. li_tunable_2021 ; zhang_room-temperature_2021 ) and Fermi surfaces for the FM phase of single layer CrTe2. The nesting vectors ${\mathbf{q}}_{M}$ and ${\mathbf{q}}_{K}$ along $\Gamma-M$ and $\Gamma-K$ directions are close to the vectors ${\mathbf{Q}}_{M}=\pi(1,1\sqrt{3})/a$ and ${\mathbf{Q}}_{K}$. By parameterizing the dispersion near the respective pairs of points of the Fermi surface by $e_{\mathbf{k}}=v_{F\sigma}\tilde{k}_{x}+\tilde{k}_{y}^{2}/(2m_{\sigma})$ where $\tilde{k}_{x,y}$ are local coordinates, assuming opposite Fermi velosities and masses of different spin states, we find at small temperatures the contribution to the exchange interactions
@@ Appendix A Single layer band structure in the ferromagnetic phase
where $\rho_{\sigma}=1/|m_{\sigma}v_{F\sigma}|$ are the curvatures of the Fermi surface sheets, $\Delta$ is the sverage spin splitting. Although the Fermi velocities are somewhat different at the respective sheets of the Fermi surfaces, this does not prevent the nesting property, as far as the difference of curvatures $\rho_{\uparrow}-\rho_{\downarrow}$ remains small close to the considered points, connected by the vectors ${\mathbf{q}}_{M}$ and ${\mathbf{q}}_{K}$.
Figure 14: (Color online) Band structure (a) and DFT density of states (b) of single-layer CrTe2 in the FM phase. Blue (red) lines correspond to majority (minority) spin states. The inset in (a) shows the DFT Fermi surface. Points show parts of the Fermi surfaces connected by nesting vectors ${\mathbf{q}}_{M}$ and ${\mathbf{q}}_{K}$; the wave vectors ${\mathbf{Q}}_{M}$ and ${\mathbf{Q}}_{K}$ are shown for comparison.
Appendix B Energy mapping
To estimate exchange interactions from the energies of various magnetic states, we perform calculations on the $2\times 2\sqrt{3}$ supercell as shown on Fig. 15. We consider 4 different spin configurations to estimate exchange parameters. For all configurations the same lattice structure was used. Each atom has 6 nearest neighbours, 6 next nearest neighbours and 6 next to next nearest neighbours and we have 8 atoms in cell. To obtain corresponding energies it is enough to consider one of the atoms and multiply the result by 8 atoms.

### v3

@@ III Results
We also used an alternative method to obtain exchange interactions, based on the comparison of energies of various spin configurations li_tunable_2021 , see Appendix B. The comparison of the obtained exchange interactions is presented in Table 1. All DFT approaches, except Ref. li_tunable_2021 yield the same signs of nearest- and next-nearest neighbor exchange interactions, with different magnitude. In particular, the nearest neighbour exchange interaction is negative (except Ref. li_tunable_2021 ) showing the tendency to antiferromagnetism. We find the sAFM-zz configuration in DFT approach 10 meV/f.u. lower than FM. At the same time, in DMFT approach we find weak ferromagnetic nearest neighbor exchange interaction, and the tendency to antiferromagnetism originates from the next to next nearest neighbors and further interactions.
Figure 7: (Color online) Exchange interaction in bilayer CrTe2 at $\beta=18$ eV-1 ($T=645$ K) (a) between the atoms of the same layer $J^{11}_{\mathbf{q}}$ and (b) between the atoms of different layers $J^{12}_{\mathbf{q}}$. Red solid line (squares) corresponds to DFT+DMFT approach in paramagnetic phase, blue dashed line (triangles) is the result of the DFT approach in FM phase. Full (open) symbols in (b) correspond to the real (imaginary) part.
@@
To study the effect of interlayer coupling we consider bilayer and trilayer CrTe2. The band structure of bilayer CrTe2 in paramagnetic phase is shown in Fig. 6. One can see that due to bilayer splitting the central sheet of the Fermi surface appears at the $\Gamma$ point. The electronic self-energies are shown in Fig. 2. One can see that the electronic damping for the bilayer system is even larger than for the monolayer one, and the non-quasiparticle form of the self-energy is also more pronounced.
@@
We consider the interaction between the atoms of the same Cr plane (intralayer part), $J^{rr}_{\mathbf{q}}$ and the interplane interaction $J^{rr^{\prime}}_{\mathbf{q}}$ with $r\neq r^{\prime}$. From the momentum dependence of intralayer exchange interactions $J^{11}_{\mathbf{q}}=J^{22}_{\mathbf{q}}$ (see Fig. 7a) one can see that an additional sheet of the Fermi surface strongly affects exchange interactions: the intralayer interaction becomes maximal near the $\Gamma$ point of the Brillouin zone, while the height of the local maximum at the $K$ point decreases in comparison to the monolayer system. The obtained maximum of the intralayer exchange interaction can be contrasted to the results of DFT approach of the present study (shown by dashed line in Fig. 7 a) and the earlier DFT study zhu_insight_2023 , where the antiferromagnetic ground state was obtained. In case of bilayer system, both the DFT and the DFT+DMFT results for exchange interactions are weakly temperature dependent because of the absence of nesting and shift of the peak of the density of states from the Fermi level. We have also verified that the tendency to the in-plane ferromagnetic order in DFT+DMFT approach in paramagnetic phase is stable for $a>3.73$Å.
Figure 10: (Color online) Real parts of the exchange interactions $J_{ij}({\mathbf{q}})\equiv J^{ij}_{\mathbf{q}}$ in trilayer system at $\beta=18$ eV-1 (645 K) in DFT+DMFT approach. The pairs of indexes $ij$ at the exchange interactions correspond to the pairs of layers (1-bottom, 2-middle, 3-top layer)
To further demonstrate the possibility of intralayer ferromagnetic order in the bilayer CrTe2, we also show the temperature dependence of the inverse uniform intralayer susceptibility $(\chi^{11,22}_{{\mathbf{q}}=0})^{-1}$ in Fig. 8. One can see that in contrast to monolayer CrTe2 the inverse intralayer susceptibility vanishes at the DMFT Curie temperature $T_{C}^{\rm DMFT,{\text{2}{\text{-}}layer}}\simeq 400K$. Similarly to the monolayer system, this DMFT Curie temperature does not determine the true Curie temperature of the system, but rather the scale of the onset of strong ferromagnetic correlations. The temperature dependence of the local magnetic susceptibility (see the inset of Fig. 8) is similar to the monolayer case, and it is characterized by the magnetic moment $\mu^{2}_{\rm loc}=33.1\mu_{B}^{2}$, which is also close to that for the monolayer compound. Therefore the strong change of the non-local properties when passing from single- to bilayer CrTe2 almost does not affect the local properties.
@@
Let us consider what changes in DFT+DMFT results when adding one more layer in trilayer CrTe2. The Fermi surface possesses on additional sheet near the $\Gamma$ point (see Fig. 9). We find that the self-energies of trilayer system (not shown) are comparable to those of the bilayer one in Fig. 2. The maximum of the intralayer exchange interaction (see Fig. 10) shifts to the $\Gamma$ point (instead of being in the near vicinity of the $\Gamma$ point in bilayer system), which shows further stabilization of intralayer ferromagnetic order with increase of the number of layers. The exchange interaction between adjacent layers is negative (i.e. antiferromagnetic), similar to the bilayer compound, with the maximum of the interlayer exchange interactions at the incommensurate positions in the $\Gamma-M$ and $\Gamma-K$ directions. We find that the tendency to incommensurate interlayer order is stable with further lowering the temperature.
@@
Figure 12: (Color online) Exchange interaction in bulk CrTe2 at $\beta=24$ eV-1 ($T=480$ K) in DFT+DMFT approach in paramagnetic phase (red solid line with circles) and the DFT approach in the FM phase (blue dashed line with filled triangles).
@@
To understand the effect of adding more layers in thick films, we consider bulk CrTe2. The obtained band structure and Fermi surface (Fig. 11) reproduce those of previous study freitas_ferromagnetism_2015 . The Fermi surface sheet closest to the $\Gamma$ point acquires the dispersion along $k_{z}$ axis in agreement with the results for the trilayer system. The exchange interactions at $\beta=24$ eV-1 ($T=480$ K, see Fig. 12) has local maximum at the $\Gamma$ point, but its value at the A point of the Brillouin zone is somewhat larger than at $\Gamma$, showing the tendency to alternating orientation between adjacent planes, similar to that obtained for 2- and 3-layer system. However, with a reduction of temperature (see Fig. 13), we observe a clear tendency of changing the global maximum of exchange interaction to $\Gamma$ point, which happens just at $T_{C}\simeq 250$ K, comparable to the experimental Curie temperature. This implies formation of low temperature FM phase (in agreement with the experimental data), which changes upon heating to a state with long- or short-range antiferromagnetic order. From the obtained exchange interactions, we find using RPA approach RPA_TC (see also Refs. MyCo ; CrO2 ), that the Neel temperature of antiferromagnetic state approximately coincides with the obtained $T_{C}$. Therefore, we expect that the long-range antiferromagnetic order above $T_{C}$ is destroyed by fluctuations. Consequently, we find an intriguing physical picture of competition of different types of magnetic correlations, which occurs just above the Curie temperature.
In view of the obtained results for the bilayer and trilayer systems, with decrease of the number of layers we expect the interlayer antiferromagnetic and incommensurate orders to become progressively more favorable. Therefore, we expect the transition to interlayer incommensurate or antiferromagnetic states with a decrease of the number of layers. In view of the delicate competition of various states, additional interactions, not included in the considered model (such as $d$-$p$ and $p$-$p$ interactions) may be important for obtaining the critical number of layers for this transition, cf. Ref. Interlayer .
@@ IV Conclusions
In summary, we have investigated magnetic properties of monolayer, bilayer, trilayer, and bulk CrTe2. In monolayer system in both the DFT and DFT + DMFT approaches, we find a preference for the incommensurate magnetic order with the wave vector ${\mathbf{Q}}_{K}$, which corresponds to 120°  spin alignment. This order competes with the AFM-zz magnetic order, previously suggested for monolayer system.
In the bilayer and trilayer CrTe2 in the DFT+DMFT approach, we find, in contrast to the DFT results, the tendency to the intralayer ferromagnetic order, which appears due to electronic correlations. At the same time, similarly to previous DFT$+U$ studies, we find a tendency to the interlayer AFM order. This tendency is preserved at not too low temperatures in bulk CrTe2, and changes to FM order at low temperatures.
The most important result of the present study is the possibility of stabilizing the intralayer FM order in CrTe2 by correlations for the number of layers greater than one. Description of the transition from FM to AFM interlayer coupling with the decrease of the number of layers requires considering more sophisticated models, including non-local $d$-$p$ (as well as $p$-$p$) interaction. The FM order in monolayer CrTe2, observed experimentally zhang_room-temperature_2021 ; meng_anomalous_2021 ; liu_wafer-scale_2023 , requires further studies, and likely appears as an effect of substrate, cf. Ref. wang_strain-_2024 .
For future studies, account of full SU(2) Coulomb interaction as well as the effect of the spin-orbit coupling would be desirable. Also, considering the effect of substrate, e.g., within the recently proposed approach of Ref. Substr is of certain interest.
@@ Acknowledgements
We are grateful to I. A. Goremykin for help with the Quantum Espresso and Wannier90 packages and discussions, and to A. K. Nukhov for discussions of the obtained results. The paramagnetic DFT+DMFT calculations and their analysis were supported by the Russian Science Foundation (Project No. 24-12-00186). The development of a program of finite frequency box corrections to the Bethe-Salpeter equation is supported within the theme “Quant” 122021000038-7 of Ministry of Science and Higher Education of the Russian Federation.
Appendix A Monolayer band structure in the ferromagnetic phase
In Fig. 14 we show the band structure (cf. Refs. li_tunable_2021 ; zhang_room-temperature_2021 ) and Fermi surfaces for the FM phase of the monolayer CrTe2. The nesting vectors ${\mathbf{q}}_{M}$ and ${\mathbf{q}}_{K}$ along the $\Gamma-M$ and $\Gamma-K$ directions are close to the vectors ${\mathbf{Q}}_{M}=\pi(1,1\sqrt{3})/a$ and ${\mathbf{Q}}_{K}$. By parameterizing the dispersion near the respective pairs of points of the Fermi surface by $e_{\mathbf{k}}=v_{F\sigma}\tilde{k}_{x}+\tilde{k}_{y}^{2}/(2m_{\sigma})$ where $\tilde{k}_{x,y}$ are local coordinates, assuming opposite Fermi velosities and masses of different spin states, we find at small temperatures the contribution to the exchange interactions
@@ Appendix A Monolayer band structure in the ferromagnetic phase
where $\rho_{\sigma}=1/(m_{\sigma}v_{F\sigma})$ are the curvatures of the Fermi surface sheets, $\Delta$ is the sverage spin splitting. Although Fermi velocities are somewhat different at the respective sheets of Fermi surfaces, this does not prevent the nesting property, as long as the difference of curvatures $\rho_{\uparrow}-\rho_{\downarrow}$ remains small close to the considered points, connected by the vectors ${\mathbf{q}}_{M}$ and ${\mathbf{q}}_{K}$.
Figure 14: (Color online) Band structure (a) and DFT density of states (b) of monolayer CrTe2 in the FM phase. Blue (red) lines correspond to majority (minority) spin states. The inset in (a) shows the DFT Fermi surface. Points show parts of the Fermi surfaces connected by nesting vectors ${\mathbf{q}}_{M}$ and ${\mathbf{q}}_{K}$; the wave vectors ${\mathbf{Q}}_{M}$ and ${\mathbf{Q}}_{K}$ are shown for comparison.
Appendix B Energy mapping of monolayer system
@@
To estimate exchange interactions from the energies of various magnetic states, we performed calculations on the $2\times 2\sqrt{3}$ supercell, as shown in Fig. 15. We consider 4 different spin configurations to estimate the exchange parameters. For all configurations the same lattice structure was used. Each atom has six nearest neighbors, six next nearest neighbours and six next to next nearest neighbours and we have eight atoms in the cell. To obtain corresponding energies it is enough to consider one of the atoms and multiply the result by 8 atoms.